# MLLM-Tool: A Multimodal Large Language Model For Tool Agent Learning

### v2

@@ Abstract
Recently, the astonishing performance of large language models (LLMs) in natural language comprehension and generation tasks triggered lots of exploration of using them as central controllers to build agent systems. Multiple studies focus on bridging the LLMs to external tools to extend the application scenarios. However, the current LLMs’ perceiving tool-use ability is limited to a single text query, which may result in ambiguity in understanding the users’ real intentions. LLMs are expected to eliminate that by perceiving the visual- or auditory-grounded instructions’ information. Therefore, in this paper, we propose MLLM-Tool, a system incorporating open-source LLMs and multi-modal encoders so that the learnt LLMs can be conscious of multi-modal input instruction and then select the function-matched tool correctly. To facilitate the evaluation of the model’s capability, we collect a dataset featured by consisting of multi-modal input tools from HuggingFace. Another important feature of our dataset is that our dataset also contains multiple potential choices for the same instruction due to the existence of identical functions and synonymous functions, which provides more potential solutions for the same query. The experiments reveal that our MLLM-Tool is capable of recommending appropriate tools for multi-modal instructions. Codes and data are available at https://github.com/MLLM-Tool/MLLM-Tool.
@@ 1 Introduction
Large Language Models (LLMs)  [2, 20, 25, 33] have showcased remarkable success in conversational scenarios. With the promising capability of the instruction following via reinforcement learning from human feedback (RLHF)  [1, 4, 13], developers extend LLMs, e.g., ChatGPT with a Function Call feature, to the tool-use domain. This extension of LLMs’ capabilities can be applied to a broad spectrum of situations, from intriguing outer tasks like restaurant reservations and online shopping to inner operations involving the retrieval and processing of infinite external knowledge. Along this direction, numerous intelligent agents  [23, 31, 24, 41, 3, 43, 7, 18] have been developed, by leveraging LLMs to accomplish various tasks in human-like ways.
@@
To remedy the previous works only accepting single text instructions, in this paper, we introduce a novel system, MLLM-Tool, integrating multi-modal encoders with open-source LLMs to synthesize multi-modal information for correct external tool identification. The distinction between our work and previous works is delineated in Table 1. Compared to previous open-sourced works, our system is the first attempt to train a large multi-modal model for tool agent learning. In light of the fact that existing tool learning benchmarks  [15, 12, 39, 23, 21, 37, 42, 27], do not explicitly construct text-to-any-modality pairs for alleviating text ambiguity, which cannot facilitate the training of our MLLM-Tool, in this paper, we collect a novel dataset, named ToolMMBench from the HuggingFace platform, one of the largest machine learning communities containing thousands of APIs. It is also worth noting that our ToolMMBench supports multi-options for each query, as demonstrated in Table 1. This feature is more closely related to real-life scenarios, where multiple alternatives exist to address the same challenges, albeit with varying performance levels. Such a feature would also make our dataset suitable for other applications in future, such as the study of integrating multiple outputs from different tools to better feedback to the user.
@@
  • We develop MLLM-Tool, a multi-modal tool agent system that can perceive the visual- or audio-grounded instructions’ information via incorporating multi-modal encoders and open-source large language models.
@@
  • We design some evaluation metrics and conduct experiments with extensive subset ablations on multiple popular LLMs, and by fine-tuning MLLM-Tool on ToolMMBench, we observe a promising accuracy, 88.19% of the tool selection. which demonstrates the effectiveness of our method.
@@ 2 Related Works
Multi-modal Large Language Models. LLMs have increasingly become the cornerstone of many natural language processing (NLP) applications. Subsequently, many efforts, including GPT-4V  [19], LLaVA  [16], BLIP-2  [14], Minigpt-4  [46] developed the LLM to vision space by aligning off-the-shelf frozen pre-trained image encoders. And similar ideas are generalized to video and audio modalities, like NeXTGPT  [38], and VideoLLaMA  [44]. Our proposed system, MLLM-Tool, focuses on exploring tool usage awareness and selection. Our goal is to empower the LLM’s ability to be conscious of the external tools’ function and select the suitable tools, to act as a controller, instead of enhancing the ability of multi-modal understanding and reasoning.
Tool Agent. Recent advancements have witnessed a rapid emergence of systems that deploy LLMs as a central orchestrator for tool assistance. Toolformer  [26] pioneers the study of bridging the LLMs with the external tools. Subsequent works have expanded the application domains of this integration to encompass health support  [17], code synthesis  [15, 30], web searching  [28]. Gorilla  [21] collected a machine learning domain dataset and introduced API retrieval to boost the performance. ToolLLM  [23] focused on executing complex tasks in practical scenarios. GPT4Tools  [42] and Visual ChatGPT  [37] integrating visual found models after decomposing the task. On the contrary, we concentrate on completely relying on the ability of large language models to naturally process multi-modal inputs and select the appropriate APIs. And to the best of our knowledge, we are the first to explore this direction.

### v3

@@ Abstract
Recently, the astonishing performance of large language models (LLMs) in natural language comprehension and generation tasks triggered lots of exploration of using them as central controllers to build agent systems. Multiple studies focus on bridging the LLMs to external tools to extend the application scenarios. However, the current LLMs’ ability to perceive tool use is limited to a single text query, which may result in ambiguity in understanding the users’ real intentions. LLMs are expected to eliminate that by perceiving the information in the visual- or auditory-grounded instructions. Therefore, in this paper, we propose MLLM-Tool, a system incorporating open-source LLMs and multi-modal encoders so that the learned LLMs can be conscious of multi-modal input instruction and then select the function-matched tool correctly. To facilitate the evaluation of the model’s capability, we collect a dataset featuring multi-modal input tools from HuggingFace. Another essential feature of our dataset is that it also contains multiple potential choices for the same instruction due to the existence of identical functions and synonymous functions, which provides more potential solutions for the same query. The experiments reveal that our MLLM-Tool is capable of recommending appropriate tools for multi-modal instructions. Codes and data are available at github.com/MLLM-Tool/MLLM-Tool.
@@ 1 Introduction
Large Language Models (LLMs)  [3, 18, 23, 31] have showcased remarkable success in conversational scenarios.
With the promising capability of the instruction following via reinforcement learning from human feedback (RLHF)  [2, 5, 11], developers extend LLMs, e.g., ChatGPT with a Function Call feature, to the tool-use domain. This extension of LLMs’ capabilities can be applied to a broad spectrum of situations, from intriguing outer tasks like restaurant reservations and online shopping to inner operations involving the retrieval and processing of infinite external knowledge. Along this path, numerous intelligent agents  [21, 29, 22, 38, 4, 40, 7, 16] have been developed, by leveraging LLMs to accomplish various tasks in human-like ways.
@@
To remedy the previous works only accepting single text instructions, in this paper, we introduce a novel system, MLLM-Tool, integrating multi-modal encoders with open-source LLMs to synthesize multi-modal information for correct external tool identification. Table 1 delineates the distinction between our work and previous works. In light of the fact that existing tool learning benchmarks  [13, 10, 36, 21, 19, 34, 39, 25], do not explicitly construct text-to-any-modality pairs for alleviating text ambiguity, which cannot facilitate the training of our MLLM-Tool, in this paper, we collect a novel dataset, named ToolMMBench from the HuggingFace platform, one of the largest machine learning communities containing thousands of APIs. It is also worth noting that our ToolMMBench supports multi-options for each query, as demonstrated in Table 1. This feature is more closely related to real-life scenarios, where multiple alternatives exist to address the same challenges, albeit with varying performance levels. Such a feature would also make our dataset suitable for other applications in the future, such as the study of integrating multiple outputs from different tools to provide better feedback to the user. The process of our MLLM-Tool is shown in Figure 2.
Figure 2: Qualitative example of MLLM-Tool. MLLM-Tool first receives the text instruction and perceives information from the input image to determine the corresponding task, selects the appropriate API that meets the requirements from the corpus, calls the API to execute, and finally gets the quality-improved image.
@@
  • We develop MLLM-Tool, a multi-modal tool agent system that can perceive the visual- or audio-grounded instructions’ information via incorporating multi-modal encoders and open-source LLMs.
@@
  • We design some evaluation metrics and conduct experiments with extensive subset ablations on multiple popular LLMs, and by fine-tuning MLLM-Tool on ToolMMBench, we observe a promising accuracy, 88.19% of the tool selection, which demonstrates the effectiveness of our method.
@@ 2 Related Works
Multi-modal Large Language Models. LLMs have increasingly become the cornerstone of many natural language processing (NLP) applications. Subsequently, many efforts include GPT-4V  [17], LLaVA  [14], BLIP2  [12], Minigpt-4  [43] developed the LLM to vision space by aligning off-the-shelf frozen pretrained image encoders. And similar ideas are generalized to video and audio modalities, like NeXTGPT  [35] and VideoLLaMA  [41]. In contrast, our proposed MLLM-Tool focuses on exploring tool usage awareness and selection. Our goal is to empower the LLM’s ability to be conscious of the external tools’ function and select the suitable tools to act as a controller instead of enhancing the multi-modal understanding and reasoning ability of MLLMs.
Tool Agent. Recent advancements have witnessed a rapid emergence of systems that deploy LLMs as a central orchestrator for tool assistance. Toolformer  [24] pioneers the study of bridging the LLMs with external tools. Subsequent works have expanded the application domains of this integration to encompass health support  [15], code synthesis  [13, 28], web searching  [26]. Gorilla  [19] collected a machine learning domain dataset and introduced API retrieval to boost the performance. ToolLLM  [21] focused on executing complex tasks in practical scenarios. GPT4Tools  [39] and Visual ChatGPT  [34] integrated visual found models after decomposing the task. It is worthwhile to note that although previous works, like HuggingGPT  [25], also handle multimodal data, the non-text data is input in a URL format. After task planning based on the text information, the selected models process the data, so the non-text data are not directly perceived. On the contrary, we concentrate on relying entirely on the ability of large language models to naturally process multimodal inputs and select the appropriate APIs, which makes the agent smart enough to understand the user’s intentions and avoid ambiguity.